OpenStudy (anonymous):

a regular pentagon has an apothem of 7.43 cm. Find the length of a side of the pentagon and the length of the radius of the circumscribed circle

OpenStudy (anonymous):

Since it is a pentagon, what are the measures of the acute angles in the right triangle formed by the apothem and the radius?

OpenStudy (anonymous):

it doesn't give me that:/

OpenStudy (anonymous):

No. You have to figure it out. First, the sum of the angles in a polygon = n-2(180) where n is the number of sides. For a pentagon, this is 5-2(180)=540. then divide by 5 for each of 5 large angles in the pentagon, and then divide by 2 for the angle formed by the radius in the right angle. This is 540/5=108 divided by 2 = 54 degrees in the right angle. Now that you know the angle , then use trig to solve the problem...

OpenStudy (anonymous):

oh! Thank you!!:D

OpenStudy (anonymous):

You are welcome!